(a) (–7) – 8 – (–25)
(b) (–13) + 32 – 8 – 1
 (d) 50 – (–40) – (–2)
 
 
smrbiombhrinietnbumtuibmsiodbm,wirjbhmriujmiosbm,

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Stdent
(a) (-7) – 8 – (-25)
= -7 -8 + 25
= -15 + 25
= 10

b) (-13) + 32 – 8 – 1
= -13 + 32 – 8 – 1
= 32 – 22
= 10

(d) 50 – (-40) – (-2)
= 50 + 40 + 2
​= 92



Thanks